# Double Check My Desired Return: Transformer with Target Alignment for Offline Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning (RL) has achieved significant advances in domains such as robotic control, autonomous driving, and medical decision-making. Most existing methods primarily focus on training policies that maximize cumulative returns from a given dataset. However, many real-world applications require precise control over policy performance levels, rather than simply pursuing the best possible return. Reinforcement learning via supervised learning (RvS) frames offline RL as a sequence modeling task, enabling the extraction of diverse policies by conditioning on different desired returns. Yet, existing RvS-based transformers, such as Decision Transformer (DT), struggle to reliably align the actual achieved returns with specified target returns, especially when interpolating within underrepresented returns or extrapolating beyond the dataset. To address this limitation, we propose Doctor, a novel approach that Double Checks the Transformer with target alignment for Offline RL. Doctor achieves superior target alignment both within and beyond the dataset, while enabling accurate and flexible control over policy performance. Notably, on the dynamic treatment regime benchmark, EpiCare, our approach effectively modulates treatment policy aggressiveness, balancing therapeutic returns against adverse event risk.
@@ 1 Introduction
In recent years, offline reinforcement learning (RL) has made remarkable progress in various domains, including robotic control (Fu et al. 2020; Levine et al. 2020), autonomous driving (Jiang et al. 2025), and medical decision-making (Hargrave, Spaeth, and Grosenick 2024). Most existing methods primarily aim to maximize the cumulative returns achieved by the learned agent, seeking to extract the best possible policy from a given offline dataset.
However, in many practical applications, it is necessary to precisely control the agent’s skill level in order to achieve different performance targets, rather than merely extracting the best-performing policy. For instance, AI clinicians need to carefully balance aggressive and conservative treatment strategies according to individual patient profiles (Rivenbark, O’Connor, and Coleman 2013), and educational AI assistants must dynamically adapt content and difficulty to match the abilities of different students (Singla et al. 2021).
Recent approaches have formulated offline RL as a sequence modeling problem analogous to large-scale language modeling, learning policies through self-supervised learning. This paradigm, known as reinforcement learning via supervised learning (RvS) (Emmons et al. 2021), predicts actions for each state in a supervised manner, conditioned on historical trajectories along with a target return prompt. By specifying the desired return as input, the learned transformer model is expected to output actions that achieve that particular return level (Chen et al. 2021; Janner, Li, and Levine 2021). Thus, RvS approaches enable the adjustment of policy by varying the target return prompt.
@@
However, transformers trained via RvS typically struggle to align the achieved return with the specified target return. As illustrated in Figure 1, Decision Transformer (DT) (Chen et al. 2021), an RvS-based method, can successfully align returns only within the range of returns that are well-supported by the dataset. DT struggles to interpolate effectively across underrepresented return regions (left side of the dashed red line), and it also fails to reliably extrapolate to higher returns beyond those observed in the training data (right side of the dashed red line). This limitation suggests that training solely with RvS is insufficient for learning policies whose actual returns closely match specified targets. Therefore, developing methods that explicitly perform target alignment is essential to accurately extract desired policies from an offline dataset, thus significantly benefiting a variety of offline RL applications. Nonetheless, existing work addressing precise target alignment in offline RL remains relatively limited.
@@
To fill this research gap, we aim to address the target alignment problem for RvS transformers, namely, precisely controlling the model to perform at different levels of expertise. We propose a novel method, Doctor, which Double Checks the Transformer with target alignment for Offline RL. At inference time, we introduce a double-check mechanism that first generates a diverse set of candidate actions by scaling outputs conditioned on sampled returns and then estimates these actions using learned value functions, selecting the action whose associated value prediction most closely aligns with the specified target. This mechanism effectively broadens the return coverage and improves alignment for underrepresented or unseen target returns. In summary, our contributions are as follows.
@@
(1) We introduce a novel double-check mechanism at inference time that significantly enhances target alignment. By generating multiple actions conditioned on sampled returns and validating these candidates using value functions, Doctor improves the model interpolation within underrepresented returns and extrapolation beyond the dataset.
(2) We demonstrate that the precise target alignment enables accurate control over the expertise of the policy, providing benefits for the medical RL. Specifically, on the dynamic treatment regime benchmark EpiCare (Hargrave, Spaeth, and Grosenick 2024), Doctor effectively modulates the aggressiveness of treatment policies, flexibly balancing therapeutic efficacy against adverse event risk.
@@
(3) We empirically evaluate Doctor on offline RL benchmarks, including D4RL (Fu et al. 2020) and EpiCare. Doctor achieves superior target alignment both within and beyond the dataset. Furthermore, in terms of return maximization, Doctor achieves competitive results compared to existing methods, showing its effectiveness in a wide range of tasks.
@@ 2 Related Work
Value Function Learning in Offline RL. Offline reinforcement learning only uses existing data collected by unknown policies without additional online data collection. One line of work is based on temporal difference (TD) learning (Zhang and Yu 2020; Fujimoto, Meger, and Precup 2019; Wu, Tucker, and Nachum 2019; Kumar et al. 2019, 2020; Sutton and Barto 2018; Dong, Ding, and Zhang 2020). To constrain the distance between the learned policy and the behavior policy to avoid distributional shift, they use a conservative value function to estimate the value of actions either by adding a regularization term in TD learning (Wu, Tucker, and Nachum 2019; Nair et al. 2020; Fujimoto and Gu 2021; Wu et al. 2022), or updating the value function in an in-sample manner (Zhou, Bajracharya, and Held 2021; Kostrikov, Nair, and Levine 2022; Zhang et al. 2023; Xiao et al. 2023). These methods aim to extract the best possible policy from the existing dataset. Unlike these approaches, Doctor’s target alignment enables the model to perform at multiple skill levels, thereby benefiting offline RL.
Sequence Modeling based methods for Offline RL. Another line of work is Reinforcement Learning via Supervised Learning (RvS). These methods cast offline RL as a conditional sequence modeling problem and learn a policy auto-regressively. These approaches (Wang, Hunt, and Zhou 2022; Lee et al. 2022; Liu and Abbeel 2023; Wu, Wang, and Hamaya 2023; Hansen-Estruch et al. 2023; Chen et al. 2023; Chebotar et al. 2023; Ma et al. 2024; Kim et al. 2024) commonly condition on goals and expect that the derived policy could be improved by feeding a high goal, but struggle to align the actual return with the desired target.
RADT (Tanaka et al. 2024) aims to address return alignment by introducing specialized aligners, which overcomes the limitations of attention allocation. RADT focuses on improving the network architecture, whereas Doctor introduces enhancements in the learning and inference methodology; thus, the two methods are orthogonal. Moreover, RADT evaluates return alignment by testing the absolute error, while Doctor empirically demonstrates that target alignment can directly benefit dynamic treatment regimes.
Recent RvS‑based approaches, such as QDT (Yamagata, Khalil, and Santos-Rodriguez 2023), QT (Hu et al. 2024), CGDT (Wang et al. 2024), and ACT (Gao et al. 2024)—also leverage Q‑values to augment transformer policies. However, these methods focus mainly on improving the model’s stitching ability to obtain the best policy from the dataset. In contrast, our primary objective is explicit target alignment, enabling an offline‑trained transformer to accurately perform at different skill levels. Moreover, unlike prior work that depends on training‑time Q‑value regularization or critic‑based prompts, the novelty of Doctor lies in the inference‑time dynamic checking mechanism.
@@ 3 Preliminaries
Offline Reinforcement Learning
Reinforcement learning (RL) (Sutton and Barto 2018) is a paradigm of agent learning via interaction. It can be modeled as a Markov Decision Process (MDP), a 5-tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},P,\gamma)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $P(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the environment dynamics, $\mathcal{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function which is bounded, $\gamma\in[0,1]$ is the discount factor. Consider the finite horizon setting, the agent interacts with the environment for $T$ steps. Denote the state, action and reward at timestep $t$ as $s_{t}$, $a_{t}$ and $r_{t}$, a trajectory is a sequence of states, actions and rewards $\tau:=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$. The return at timestep $t$ is defined as $R_{t}=\sum_{i=t}^{T}\gamma^{i-t}r_{i}$. The goal of an reinforcement learning agent is to learn an optimal policy $\pi$ that maximizes the expected return $R_{0}=\mathbb{E}_{\pi}[\sum_{i=0}^{T}\gamma^{i}r_{i}]$.
In offline RL, instead of interacting with the environment, the agent learns from a static dataset of trajectories $\mathcal{D}:=\{\tau_{j}\}$ such as the D4RL benchmark (Fu et al. 2020). Several recent offline‑RL methods impose an in‑sample constraint (Levine et al. 2020), learning Q-functions over the dataset’s state–action visitation:
@@
For example, Implicit Q‑Learning (IQL) (Kostrikov, Nair, and Levine 2022) only bootstraps from actions seen in the dataset, avoiding value estimates on unseen actions and substantially reducing estimation error. This setting eliminates the need for online exploration, which is practical in scenarios where exploration is expensive or dangerous.
Reinforcement Learning via Supervised Learning.
Decision Transformer (DT) applies transformers to offline RL. Different from offline RL methods based on temporal difference learning, DT models offline RL as a sequence modeling problem and learns a policy auto-regressively by predicting the next action given the history trajectory conditioned on a target return. This set of approaches abstract offline RL as a sequence modeling problem and learns a policy via supervised learning (RvS). These approaches (Chen et al. 2021; Janner, Li, and Levine 2021) frame offline RL as a sequence modeling problem. The K-step historical trajectory sequence $\tau_{t}$ consists of three modalities:
@@ 4 Method
Model Architecture
Our model adopts an encoder-decoder architecture as a universal representation extractor. Both the encoder and decoder are bidirectional transformers, which are adept at capturing dependencies in sequential data. The task is based on sequence reconstruction from masked views (He et al. 2022; Vaswani 2017; Kitaev, Kaiser, and Levskaya 2020), where a random subset of the sequence is masked and the model is tasked with reconstructing the original trajectory.
This design choice is motivated by recent work (Wu et al. 2023) which demonstrated that bidirectional transformers trained with random and auto-regressive masking encourage the model to reconstruct missing tokens by leveraging the context of the unmasked sequence, enhancing the generalization. We expect this design to yield stronger interpolation when handling trajectories associated with underrepresented returns. Our experiments (Figure 4 and Appendix E.2) confirm that the bidirectional transformer–based architecture achieves improved target alignment.
@@
Training
Our training consists of two purposes: (1) reconstructing the original trajectory sequence from the masked input trajectory, and (2) learning the action-value $q_{t}$ to double check the candidates to improve alignment.
@@
$\mathcal{L}_{\text{recon}}(\theta)=-\sum_{t=0}^{T}\Bigl{(}\log P_{\theta}(R_{t}|M(\tau_{t}))+\\ \log P_{\theta}(s_{t}|M(\tau_{t}))+\log P_{\theta}(a_{t}|M(\tau_{t}))\Bigr{)}.$ (4)
@@
The goal of the $Q$ function is to learn an optimal action-value function, aligning with target returns. To avoid querying the learned $Q$ function on out-of-sample actions, we utilize the asymmetric least squares loss function (Kostrikov, Nair, and Levine 2022). Denote the learnable parameters of the $Q$ function as $\phi$. The learning loss is defined by:
@@
where $r_{t}$ is the reward, $\gamma$ is the discount factor, $Q_{\phi,t}$ and $Q_{\phi,t+1}$ are the $Q$ functions at time step $t$ and $t+1$, respectively. $L_{2}^{\nu}(u)=\lvert\nu-\mathbf{1}(u<0)\rvert u^{2}$ is the asymmetric least squares loss function. For $\nu=0.5$, the loss function is equivalent to the standard mean squared error loss. For $\nu>0.5$, the loss function is asymmetric, which down-weights the contributions of values smaller than zero (Newey and Powell 1987; Kostrikov, Nair, and Levine 2022). Eq. (1) can be approximated with $\tau\approx 1$.
@@
Inference Time Alignment
Given a tuple $(R_{t},s_{t},a_{t},q_{t})$, the return $R_{t}$ represents the return in the dataset when taking action $a_{t}$ at state $s_{t}$, while the action-value $q_{t}$ reflects the expected return. We expect there could be a gap between $q_{t}$ and $R_{t}$. This gap presents the difference between policy evaluation of the unknown behavior policy that collected the dataset and the best possible policy after policy improvement. This motivates us to introduce a double-check mechanism during inference to ensure alignment between the predicted action and the target return.

### v2

@@ Abstract
Offline reinforcement learning (RL) has achieved significant advances in domains such as robotic control, autonomous driving, and medical decision-making. Most existing methods primarily focus on training policies that maximize cumulative returns from a given dataset. However, many real-world applications require precise control over policy performance levels, rather than simply pursuing the best possible return. Reinforcement learning via supervised learning (RvS) frames offline RL as a sequence modeling task, enabling the extraction of diverse policies by conditioning on different desired returns. Yet, existing RvS-based transformers, such as Decision Transformer (DT), struggle to reliably align the actual achieved returns with specified target returns, especially when interpolating within underrepresented returns or extrapolating beyond the dataset. To address this limitation, we propose Doctor, a novel approach that Double Checks the Transformer with target alignment for Offline RL. Doctor integrates the strengths of supervised learning (SL) and temporal difference (TD) learning by jointly optimizing the action prediction and value estimation. During inference, Doctor introduces a double-check mechanism: actions are first sampled around the desired target returns and then validated with value functions. This ensures more accurate alignment between predicted actions and desired target returns. We evaluate Doctor on the D4RL and EpiCare benchmarks, demonstrating aligned control yields stronger performance and tunable expertise, showing its effectiveness in a wide range of tasks.
keywords:
Offline Reinforcement Learning, Transformer, Target Alignment, Decision and Control
journal: arXiv
[…]
[label1]organization=School of Artificial Intelligence, Beihang University, city=Beijing, country=China
[…]
[label2]organization=Department of Computing Science and Amii, University of Alberta, city=Edmonton, country=Canada
[…]
[label3]organization=Edmonton Research Center, Huawei Canada, city=Edmonton, country=Canada
[…]
[label4]organization=School of Artificial Intelligence and Computer Science, North China University of Technology, city=Beijing, country=China
[…]
[label5]organization=School of Computer Science and Engineering, Beihang University, city=Beijing, country=China
[…]
[label6]organization=Research Institute of Multiple Agents and Embodied Intelligence, Peng Cheng Laboratory, city=Shenzhen, country=China
@@ 1 Introduction
In recent years, offline reinforcement learning (RL) [1] has made remarkable progress in various domains, including robotic control [2, 3], autonomous driving [4], and medical decision-making [5, 6]. Most existing methods primarily aim to maximize the cumulative rewards, seeking to extract the best possible policy from a given offline dataset [7].
However, in many practical applications, it is necessary to precisely control the agent’s skill level in order to achieve different performance targets, rather than merely extracting the best-performing policy. For instance, AI clinicians need to satisfy personalized medicine requirements according to individual patient profiles [8], and educational AI assistants must dynamically adapt content and difficulty to match the abilities of different students [9, 10]. In game AI, diverse NPC skill levels are crucial for a rich and engaging gameplay experience [11].
Recent approaches have formulated offline RL as a sequence modeling problem analogous to large-scale language modeling, learning policies through supervised learning [12, 13]. This paradigm, known as reinforcement learning via supervised learning (RvS) [14], predicts actions for each state in a supervised manner, conditioned on historical trajectories along with a target return prompt. By specifying the desired return as input, the learned transformer model is expected to output actions that achieve that particular return level. Thus, RvS approaches enable the adjustment of policy by varying the target return prompt.
@@
However, transformers trained via RvS typically struggle to align the achieved return with the specified target return. As illustrated in Figure 1, Decision Transformer (DT) [12], an RvS-based method, only achieves alignment within well-supported returns in the dataset. DT struggles to interpolate effectively across underrepresented return regions (left side of the dashed red line), and it also fails to reliably extrapolate to higher returns beyond those observed in the training data (right side of the dashed red line). This limitation suggests that training solely with RvS is insufficient for learning policies whose actual returns closely match specified targets. Therefore, developing methods that explicitly perform target alignment is essential to accurately extract desired policies from an offline dataset, thus significantly benefiting a variety of applications. Nonetheless, existing work addressing precise target alignment in offline RL remains relatively limited.
@@
To fill this research gap, we aim to address the target alignment problem for RvS transformers, namely, achieving aligned control that calibrates policy behavior across expertise levels. We propose a novel method, Doctor, which Double Checks the Transformer with target alignment for Offline RL. Doctor integrates the strengths of supervised learning and temporal difference learning by jointly optimizing the action prediction and value estimation. Supervised learning stabilizes the prediction of actions conditioned on target returns, while temporal difference learning adds stitching capability to the transformer and plays a critical role in alignment. At inference time, we introduce a double-check mechanism that first generates a diverse set of candidate actions by scaling outputs conditioned on sampled returns and then estimates these actions using learned value functions, selecting the action whose associated value prediction most closely aligns with the specified target. This mechanism effectively broadens the return coverage and improves alignment for underrepresented or unseen target returns. In summary, our contributions are as follows.
(1) We propose a novel method, Doctor, that integrates the strengths of supervised learning and temporal difference learning in a transformer for offline RL. By jointly optimizing supervised trajectory reconstruction and temporal-difference value estimation, Doctor demonstrates the ability to interpolate across underrepresented return ranges and extrapolate beyond dataset-supported regions.
@@
(2) Doctor introduces a novel double-check mechanism at inference time that significantly enhances target alignment. By generating multiple actions conditioned on sampled returns and validating these candidates using value functions, this mechanism enhances the alignment between predicted actions and desired targets, enabling the extraction of policies with varying performance levels
@@
(3) We empirically evaluate Doctor on offline RL benchmarks, including D4RL [15] and EpiCare [5], demonstrating superior target alignment both within and beyond the dataset. Furthermore, aligned control can translate to better return maximization performance. Doctor achieves competitive results compared to existing methods, showing its effectiveness in a wide range of tasks.
(4) We demonstrate that the precise target alignment enables accurate control over the expertise of the policy, providing benefits for RL applications. Specifically, on the dynamic treatment regime benchmark EpiCare, Doctor effectively modulates the aggressiveness of treatment policies, flexibly balancing therapeutic efficacy against adverse event risk.
@@ 2 Related Work
Value Function Learning in Offline RL. Offline reinforcement learning only uses existing data collected by unknown policies without additional online data collection. One line of work is based on temporal difference (TD) learning [16, 17, 18, 19]. To constrain the distance between the learned policy and the behavior policy to avoid distributional shift, they use a conservative value function to estimate the value of actions either by adding a regularization term in TD learning [20, 21], or updating the value function in an in-sample manner [18, 22]. CQL [23] augments the standard Bellman error objective with a simple Q-value regularizer, such that the expected value of a policy under the learned Q-function lower-bounds its true value. Implicit $Q$-Learning (IQL) [18] estimates the value of the best available action at a given state with expectile regression, without ever directly querying the $Q$ function for unseen actions. These methods aim to extract the best possible policy from the existing dataset. Unlike these approaches, Doctor’s target alignment enables the model to perform at multiple skill levels, thereby benefiting offline RL.
Sequence Modeling based Methods for Offline RL. Another line of work is Reinforcement Learning via Supervised Learning (RvS). These methods cast offline RL as a conditional sequence modeling problem and learn a policy auto-regressively. Benefit from the inherent stability and scalability of SL, these methods bypasses the need for bootstrapping for long term credit assignment and avoids the “deadly triad” [17] known to destabilize RL. These approaches [24, 25, 26] commonly condition on goals and expect that the derived policy could be improved by feeding a high goal. DT [12] trains a transformer to autoregressively predict action sequences based on desired return and past trajectory. MTM [27] applies masked prediction to learn a generic and versatile model for prediction, representation, and control. To enhance transformers with stitching ability and recover the best policy from the dataset, recent approaches leverage Dynamic Programming [28, 29], exploit Q-values to guide actions [30], or incorporate Q-value regularization into transformers [31]. Although RvS tends to be stable, these methods struggle to align the actual return with the desired target return [32]. Therefore, our primary objective is explicit target alignment, enabling an offline‑trained transformer to accurately perform at different skill levels.
@@ 3 Preliminaries
3.1 Offline Reinforcement Learning
Reinforcement learning (RL) [17] is a paradigm of agent learning via interaction. It can be modeled as a Markov Decision Process (MDP), a 5-tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},P,\gamma)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $P(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the environment dynamics, $\mathcal{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function which is bounded, $\gamma\in[0,1]$ is the discount factor. Consider the finite horizon setting, the agent interacts with the environment for $T$ steps. Denote the state, action and reward at timestep $t$ as $s_{t}$, $a_{t}$ and $r_{t}$, a trajectory is a sequence of states, actions and rewards $\tau:=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$. The return at timestep $t$ is defined as $R_{t}=\sum_{i=t}^{T}\gamma^{i-t}r_{i}$. The goal of an reinforcement learning agent is to learn an optimal policy $\pi$ that maximizes the expected return $R_{0}=\mathbb{E}_{\pi}[\sum_{i=0}^{T}\gamma^{i}r_{i}]$.
In offline RL, instead of interacting with the environment, the agent learns from a static dataset of trajectories $\mathcal{D}:=\{\tau_{j}\}$ such as the D4RL benchmark [15]. The dataset is collected by either one or multiple unknown behavior policies, and the agent’s goal is to learn a policy based on the dataset that performs well in the environment. Several recent offline‑RL methods impose an in‑sample constraint, learning Q-functions over the dataset’s state–action visitation:
@@ 3.1 Offline Reinforcement Learning
For example, Implicit Q‑Learning (IQL) [18] only bootstraps from actions seen in the dataset, avoiding value estimates on unseen actions and substantially reducing estimation error. This setting eliminates the need for online exploration, which is practical in scenarios where exploration is expensive or dangerous.
3.2 Reinforcement Learning via Supervised Learning.
The transformer model [33] is trained with a masked language modeling objective, where the model predicts the next token in the sequence given the previous tokens. Transformers have shown remarkable success in various tasks such as natural language processing and computer vision [34, 35]
Decision Transformer (DT) [12] applies transformers to offline RL. Different from offline RL methods based on temporal difference learning, DT models offline RL as a sequence modeling problem and learns a policy auto-regressively by predicting the next action given the history trajectory conditioned on a target return. This set of approaches abstract offline RL as a sequence modeling problem and learns a policy via supervised learning (RvS). The K-step historical trajectory sequence $\tau_{t}$ consists of three modalities:
@@ 4 Method
4.1 Model Architecture
Our model adopts an encoder-decoder architecture as a universal representation extractor. Both the encoder and decoder are bidirectional transformers, which are adept at capturing dependencies in sequential data. The task is based on sequence reconstruction from masked views [33, 36], where a random subset of the sequence is masked and the model is tasked with reconstructing the original trajectory.
This design choice is motivated by recent work [27] which demonstrated that bidirectional transformers trained with random and auto-regressive masking encourage the model to reconstruct missing tokens by leveraging the context of the unmasked sequence, enhancing the generalization. We expect this design to yield stronger interpolation when handling trajectories associated with underrepresented returns. Our experiments (Figure 4) confirm that the bidirectional transformer–based architecture achieves improved target alignment.
@@ 4.1 Model Architecture
4.2 Training
Our training consists of two purposes: (1) reconstructing the original trajectory sequence from the masked input trajectory, which is a self-supervised learning task, and (2) learning the action-value $q_{t}$ to enable the model for stitching and to improve the alignment, which is TD learning. We jointly optimize the model to minimize the reconstruction error and the TD error.
@@ 4.2 Training
$\mathcal{L}_{\text{recon}}(\theta)=-\sum_{t=0}^{T}\Bigl(\log P_{\theta}(R_{t}|M(\tau_{t}))+\log P_{\theta}(s_{t}|M(\tau_{t}))+\log P_{\theta}(a_{t}|M(\tau_{t}))\Bigr).$ (4)
@@
The goal of the $Q$ function is to learn an optimal action-value function, aligning with target returns. To avoid querying the learned $Q$ function on out-of-sample actions, we utilize the asymmetric least squares loss function [18]. Denote the learnable parameters of the $Q$ function as $\phi$. The learning loss is defined by:
@@
where $r_{t}$ is the reward, $\gamma$ is the discount factor, $Q_{\phi,t}$ and $Q_{\phi,t+1}$ are the $Q$ functions at time step $t$ and $t+1$, respectively. $L_{2}^{\nu}(u)=\lvert\nu-\mathbf{1}(u<0)\rvert u^{2}$ is the asymmetric least squares loss function. For $\nu=0.5$, the loss function is equivalent to the standard mean squared error loss. For $\nu>0.5$, the loss function is asymmetric, which down-weights the contributions of values smaller than zero [37, 18, 38]. Eq. (1) can be approximated with $\tau\approx 1$.
@@
4.3 Inference Time Alignment
Given a tuple $(R_{t},s_{t},a_{t},q_{t})$, the return $R_{t}$ represents the return in the dataset when taking action $a_{t}$ at state $s_{t}$, while the action-value $q_{t}$ reflects the expected return.
We expect there could be a gap between $q_{t}$ and $R_{t}$. This gap presents the difference between policy evaluation of the unknown behavior policy that collected the dataset and the best possible policy after policy improvement. This motivates us to introduce a double-check mechanism during inference to ensure alignment between the predicted action and the target return.